OpenStudy (anonymous):

Solve the inequality Write the solution in set builder notation -2x-13≥-25 ? HELP PLEASE !

OpenStudy (anonymous):

what is a set builder notation? anyways i can help you solve the inequality it works mostly like if you were to solve an equation so -2x-13 (>or=) -25 sorry i cant type symbols add 13 to both sides

OpenStudy (anonymous):

then you get -2x (> or =) -12

OpenStudy (anonymous):

now to get x on its own you divide by -2 from both sides

OpenStudy (kamille):

\[-2x \ge-25+13\] add 25+13 and then divide by (-2). Don't forget to change sign into another (from \[from \ge \to l\]

OpenStudy (kamille):

\[from \ge \] to <(or=)

OpenStudy (anonymous):

I just need the answer because I don't really get nothing of this.

OpenStudy (anonymous):

now you are not done yet whenever you have to divide or multiply by a negetive number you have to change the sign so the answer is NOT: x (> or =) 6 but it is: x (< or =) 6

OpenStudy (anonymous):

did you get it now @Biancao9o3o12 ?

OpenStudy (anonymous):

Thank you Very much Kamille and Bhagyashree

OpenStudy (anonymous):

you welcome!